OpenStudy (anonymous):

Three CDs are picked at random from a collection of 13 CDs of which six are defective. Find the probability of the following events. (Round your answers to three decimal places.) (a) All three CDs are defective. (b) All three CDs are functioning properly.

jimthompson5910 (jim_thompson5910):

How many ways are there to select 3 defective CDs?

OpenStudy (anonymous):

@jim_thompson5910 yes

jimthompson5910 (jim_thompson5910):

I'm asking for a number. Re-read what I asked above.

OpenStudy (anonymous):

I was thinking (13!)/(10!) but that just seems off

jimthompson5910 (jim_thompson5910):

there are 6 defective CDs and 3 slots n = 6 r = 3 n C r = (n!)/(r!*(n-r)!) 6 C 3 = (6!)/(3!*(6-3)!) 6 C 3 = ???

jimthompson5910 (jim_thompson5910):

I'm using n C r instead of n P r because order doesn't matter

OpenStudy (anonymous):

So if we do the 6 C 3 I receive the value of 20

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

how many ways are there to select 3 CDs (defective or nondefective, it doesn't matter)

OpenStudy (anonymous):

So n = the total amount and r = the possible ways ?

jimthompson5910 (jim_thompson5910):

yeah if you had 5 people and 3 seats, then n = 5 and r = 3

jimthompson5910 (jim_thompson5910):

how many ways are there to select 3 CDs (defective or nondefective, it doesn't matter) ?

OpenStudy (anonymous):

13 = n

OpenStudy (anonymous):

then 6-3 = r?

jimthompson5910 (jim_thompson5910):

r = number of slots

jimthompson5910 (jim_thompson5910):

r = 3 slots

OpenStudy (anonymous):

I use the same formula correct?

jimthompson5910 (jim_thompson5910):

yes use the n C r formula

OpenStudy (anonymous):

3 C 6 or 3 C 13 ( I apolgize as I am trying to learn :( )

OpenStudy (anonymous):

apologize *

jimthompson5910 (jim_thompson5910):

n = 13 r = 3 n C r = 13 C 3 = ??

jimthompson5910 (jim_thompson5910):

Rule: the value of r cannot be larger than the value of n

OpenStudy (anonymous):

286 ... so again n would = 13 because thats the total of CDS and because we want 3 CDS and it doesn't matte if there defective or not. That is how we get to 13 C 3

jimthompson5910 (jim_thompson5910):

yes, there are 286 ways to pick 3 CDs from a pool of 13 (order doesn't matter)

jimthompson5910 (jim_thompson5910):

So the probability of picking 3 defective is 20/286 = 10/143

jimthompson5910 (jim_thompson5910):

try out part b and tell me what you get

OpenStudy (anonymous):

i am sorry where did you receive the 20? from 20/286

OpenStudy (anonymous):

I will try part B thanks!

jimthompson5910 (jim_thompson5910):

from n C r = 6 C 3 = 20

OpenStudy (anonymous):

Okay so if 6 are defective then they're 7 good ones

OpenStudy (anonymous):

so would it be 7 C 3

OpenStudy (anonymous):

then that answer = 35

jimthompson5910 (jim_thompson5910):

so far, so good

OpenStudy (anonymous):

we don't use 286 right it wouldn't be 35/286

OpenStudy (anonymous):

Wait yeah. we use 286 because it is still 13 C 3

OpenStudy (anonymous):

Yes !! it is ! omg I get it now. so C = choice ...so n will always = the amount that you have then r will be what you are looking for however like you said r will never be bigger than N

OpenStudy (anonymous):

@jim_thompson5910 you are a life saver !!

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

C stands for combination to be more precise compared to the P for permutation in n P r http://www.regentsprep.org/regents/math/algtrig/ats5/lcomb.htm

OpenStudy (anonymous):

@jim_thompson5910 thank you I will check out that link YOU ARE the BEST

jimthompson5910 (jim_thompson5910):

I'm glad I could help out

OpenStudy (anonymous):

@jim_thompson5910 I wil try to use this for the rest of my questions. thanks again!